you roll a fair die 20 times and end up with getting 3 ones 4 twos 5 threes 0 fours 2 fives and 6 sixes what is the theoretical P(1,4)
1
Expert's answer
2012-11-28T10:16:21-0500
P = number offavorable outcomes / number of possible outcome
P (1) = 3/20 = 0.15 or 15%
P (4) = 0/20 = 0